# Construction of asymptotic quantum many-body scar states in the SU($N$) Hubbard model

### v2

@@ Abstract
We construct asymptotic quantum many-body scars (AQMBS) in one-dimensional SU($N$) Hubbard chains ($N\geq 3$) by embedding the scar subspace into an auxiliary Hilbert subspace $\mathcal{H}_{P}$ and identifying a parent Hamiltonian within it, together with a corresponding extension of the restricted spectrum-generating algebra to the multi-ladder case. Unlike previous applications of the parent-Hamiltonian scheme, we show that the parent Hamiltonian becomes the SU($N$) ferromagnetic Heisenberg model rather than the spin-1/2 case, so that its gapless magnons realize explicit AQMBS of the original model. Working in the doublon–holon subspace, we derive this mapping, obtain the one-magnon dispersion for periodic and open boundaries, and prove (i) orthogonality to the tower of scar states, (ii) vanishing energy variance in the thermodynamic limit, and (iii) subvolume entanglement entropy with rigorous MPS/MPO bounds. Our results broaden the parent-Hamiltonian family for AQMBS beyond spin-1/2 and provide analytic, low-entanglement excitations in SU($N$)-symmetric systems.
@@ I Introduction
AQMBS states have been constructed in various models, including the spin-1 XY model [16, 30, 32], the extended Fermi–Hubbard model with correlated hopping terms [23, 24, 33], the Heisenberg model with Dzyaloshinskii-Moriya interaction [34], the domain-wall conserving model [35, 34], the DH model [36, 37], the PXP model [12, 13, 34], the Onsager scar model [17, 38, 34], the Affleck-Kennedy-Lieb-Tasaki model [34, 39], and others [40, 41].
@@
The rest of this paper is organized as follows: In Sec. II, we review the symmetry- and algebra-based framework underlying QMBS and its extension to AQMBS, fix notation, and present the multi-ladder generalization relevant to our setting. In Sec. III, we apply this construction to the $\mathrm{SU}(N)$ Hubbard model: after introducing the model and the ladder operators, we derive the parent Hamiltonian and show that it is the $\mathrm{SU}(N)$ ferromagnetic Heisenberg chain. We then analyze the resulting AQMBS excitations, including their orthogonality to the tower of scars, vanishing energy variance, dispersion, and entanglement scaling. Section IV summarizes our findings and outlines open directions.
@@ II.1.1 Review of single ladder RSGA
Let the Hamiltonian $\hat{H}$, the operator $\hat{Q}^{\dagger}$, and the state $|S_{0}\rangle$ satisfy the following conditins, with $\hat{Q}^{\dagger}|S_{0}\rangle\neq 0$:
@@ II.1.2 Multi-ladder extension of RSGA
Under the MLRSGA-$m$ conditions, the action of each ladder operator $\{\hat{Q}_{l}^{\dagger}\}_{l=1}^{N}$ on the tower of states generated from $|S_{\bm{0}}\rangle$ includes an energy shift $\mathcal{E}_{l}$. As a result, any state obtained by applying a total of $M$ raising operations has its energy shifted by $\sum_{l}m_{l}\mathcal{E}_{l}$, where $m_{l}$ denotes the number of times the operator $\hat{Q}_{l}^{\dagger}$ is applied. More precisely, for the multi-index $\bm{m}=(m_{1},\ldots,m_{N})$, we define the unnormalized tower of states and total excitation number $M$:
$|\tilde{S}_{\bm{m}}\rangle$ $\coloneq\begin{cases}\vskip 3.99994pt\displaystyle{\prod_{l=1}^{N}(\hat{Q}_{l}^{\dagger})^{m_{l}}|S_{\bm{0}}\rangle},\quad&\text{for }M\leq M_{\rm max},\\ 0,\quad&\text{for }M>M_{\rm max},\end{cases}$ (20)
$M$ $\coloneq\sum_{l=1}^{N}m_{l},$ (21)
where $M_{\rm max}$ denotes the maximal total excitation allowed by the model. Using Eqs. (14)-(19), one obtains
@@
Thus the states $|S_{\bm{m}}\rangle$ constitute a tower of QMBS states.
@@ II.2 Symmetry-based formalism
To ensure that the the presence of a noncommuting term $\hat{C}$, we show that $\hat{C},\hat{Q}_{\bm{\alpha}}]$ is itself a ladder operator with the same root $\bm{\alpha}$. For $X=\hat{Q}_{\mu}^{z},Y=\hat{C}$, and $Z=\hat{Q}_{\bm{\alpha}}$, the Jacobi identity $[X,[Y,Z]]+[Y,[Z,Z]]+[Z,[X,Y]]=0$ gives
@@ III Results: Application to SU($N$) Hubbard Model
Previous applications of the parent-Hamiltonian scheme have yielded the spin-$1/2$ ferromagnetic Heisenberg model [33]. In what follows, we apply the formalism developed in the previous section to the SU($N$) Hubbard model, which exhibits higher internal symmetry and supports multiple ladder operators. We show that the parent Hamiltonian in this setting is the ferromagnetic SU($N$) Heisenberg model, and the AQMBS states are identified with gapless excited states of this parent Hamiltonian.
@@ III.1 Model
$\hat{H}_{\mathrm{int}}$ $\coloneq\frac{1}{2}U\sum_{j}\sum_{n,m=1,n\not=m}^{N}\hat{n}_{j,\alpha_{m}}\hat{n}_{j,\alpha_{n}},$ (59)
@@ III.2 Parent Hamiltonian
By construction, $\hat{\mathcal{P}}_{j,j+1}^{\dagger}=\hat{\mathcal{P}}_{j,j+1}$ and $\hat{\mathcal{P}}_{j,j+1}^{2}=\hat{\mathcal{P}}_{j,j+1}$. The parent Hamiltonian can be written as
@@
$=2J^{2}\left(|\alpha_{1}\alpha_{\mu},{\rm vac}\rangle-|{\rm vac},\alpha_{1}\alpha_{\mu}\rangle\right),$ (84)
@@ III.3 Gapless excitations of the parent Hamiltonian
Having established that the parent Hamiltonian is equivalent to the SU($N$) ferromagnetic Heisenberg model, we now turn to its low-energy excitations. It is well known that the ferromagnetic Heisenberg model supports gapless magnon excitations [58], which naturally provide candidates for the AQMBS states in our framework. In the following, we first impose periodic boundary conditions and construct one-magnon excitations with well-defined momentum. Subsequently, we analyze the case of open boundary conditions, where the corresponding low-energy excitations are obtained by constructing standing-wave–like states. Similar calculations have been performed in Appendix C of Ref. [33].
@@
with $p=1,2,\ldots,L$ and $\mathcal{N}_{p}$ a normalization constant. Using these eigenmodes, we define a standing-wave creation operator
@@
so that $\hat{F}^{\mu 0}_{\mathrm{OBC}}(p)|\bm{0}\rangle$ is a normalized one-magnon state. Direact calculations show that
@@
Using the commutation relation $\hat{H}_{\rm p},\hat{F}_{\rm tot}^{\mu 0}]=0$ and the fact that $|\bm{0}\rangle$ is a ground state of the SU($N$) ferromagnetic Heisenberg model, we can show that the following states are also ground states fo the parent Hamiltonian:

### v3

@@ Abstract
We construct asymptotic quantum many-body scars (AQMBS) in one-dimensional SU($N$) Hubbard chains ($N\geq 3$) by embedding the scar subspace into an auxiliary Hilbert subspace $\mathcal{H}_{P}$ and identifying a parent Hamiltonian within it, together with a corresponding extension of the restricted spectrum-generating algebra to the multi-ladder case. Unlike previous applications of the parent-Hamiltonian scheme, we show that the parent Hamiltonian becomes the SU($N$) ferromagnetic Heisenberg model rather than the spin-1/2 case, so that its gapless magnons realize explicit AQMBS of the original model. Working in the doublon–holon subspace, we derive this mapping, obtain the one-magnon dispersion for periodic and open boundaries, and prove (i) orthogonality to the scar states, (ii) vanishing energy variance in the thermodynamic limit, and (iii) subvolume entanglement entropy with rigorous MPS/MPO bounds. Our results broaden the parent-Hamiltonian family for AQMBS beyond spin-1/2 and provide analytic, low-entanglement excitations in SU($N$)-symmetric systems.
@@ I Introduction
AQMBS states have been constructed in various models, including the spin-1 XY model [16, 30, 32], the extended Fermi–Hubbard model with the correlated hopping terms [23, 24, 33], the Heisenberg model with Dzyaloshinskii-Moriya interaction [34], the domain-wall conserving model [35, 34], the DH model [36, 37], the PXP model [12, 13, 34], the Onsager scar model [17, 38, 34], the Affleck-Kennedy-Lieb-Tasaki model [34, 39], and others [40, 41].
@@
The rest of this paper is organized as follows: In Sec. II, we review the symmetry- and algebra-based framework underlying QMBS and its extension to AQMBS, fix notation, and present the multi-ladder generalization relevant to our setting. In Sec. III, we apply this construction to the $\mathrm{SU}(N)$ Hubbard model: after introducing the model and the ladder operators, we derive the parent Hamiltonian and show that it is the $\mathrm{SU}(N)$ ferromagnetic Heisenberg chain. We then analyze the resulting AQMBS excitations, including their orthogonality to scars states, vanishing energy variance, dispersion, and entanglement scaling. Section IV summarizes our findings and outlines open directions.
@@ II.1.1 Review of single ladder RSGA
Let the Hamiltonian $\hat{H}$, the operator $\hat{Q}^{\dagger}$, and the state $|S_{0}\rangle$ satisfy the following conditions, with $\hat{Q}^{\dagger}|S_{0}\rangle\neq 0$:
@@ II.1.2 Multi-ladder extension of RSGA
Under the MLRSGA-$m$ conditions, the action of each ladder operator $\{\hat{Q}_{l}^{\dagger}\}_{l=1}^{N}$ on the scar states generated from $|S_{\bm{0}}\rangle$ includes an energy shift $\mathcal{E}_{l}$. More precisely, for the multi-index $\bm{m}=(m_{1},\ldots,m_{N})$, we define the unnormalized scar states:
$|\tilde{S}_{\bm{m}}\rangle$ $\coloneq\vskip 3.99994pt\displaystyle{\prod_{l=1}^{N}(\hat{Q}_{l}^{\dagger})^{m_{l}}|S_{\bm{0}}\rangle},$ (20)
where we consider the case $|\tilde{S}_{\bm{m}}\rangle\neq 0$. Using Eqs. (14)-(19), one obtains
@@
Thus the states $|S_{\bm{m}}\rangle$ constitute a set of QMBS states.
@@ II.2 Symmetry-based formalism
To ensure that the the presence of a noncommuting term $\hat{C}$, we show that $\hat{C},\hat{Q}_{\bm{\alpha}}]$ is itself a ladder operator with the same root $\bm{\alpha}$. For $X=\hat{Q}_{\mu}^{z},Y=\hat{C}$, and $Z=\hat{Q}_{\bm{\alpha}}$, the Jacobi identity $[X,[Y,Z]]+[Y,[Z,X]]+[Z,[X,Y]]=0$ gives
@@ II.3 Systematic construction of AQMBS: review and multi-ladder extension
As shown in Ref. [33], the parent Hamiltonian reduces to the spin-$1/2$ ferromagnetic Heisenberg model in many cases. In such cases, the low-energy gapless excited states (i.e., magnon-excited states) can be obtained analytically from simple algebraic relations. However, the systematic construction does not guarantee that the parent Hamiltonian always reduces to a pure ferromagnetic Heisenberg model. When this is not the case, we need to rely on approximations or numerical calculations to obtain the AQMBS.
@@ III Results: Application to SU($N$) Hubbard Model
Previous applications of the parent-Hamiltonian scheme have yielded the spin-$1/2$ ferromagnetic Heisenberg model as described in the previous section [33]. In what follows, we apply the formalism developed in the previous section to the SU($N$) Hubbard model, which exhibits higher internal symmetry and supports multiple ladder operators. We show that the parent Hamiltonian in this setting is the ferromagnetic SU($N$) Heisenberg model, and the AQMBS states are identified with gapless excited states of this parent Hamiltonian.
@@ III.1 Model
$\hat{H}_{\mathrm{int}}$ $\coloneq\frac{U}{2}\sum_{j}\sum_{n,m=1,n\not=m}^{N}\hat{n}_{j,\alpha_{m}}\hat{n}_{j,\alpha_{n}},$ (58)
@@ III.2 Parent Hamiltonian
By construction, $\hat{\mathcal{P}}_{j,j+1}^{\dagger}=\hat{\mathcal{P}}_{j,j+1}$ and $\hat{\mathcal{P}}_{j,j+1}^{2}=\hat{\mathcal{P}}_{j,j+1}$ hold. The parent Hamiltonian can be written as
@@
$=+2J^{2}\left(|\alpha_{1}\alpha_{\mu},{\rm vac}\rangle-|{\rm vac},\alpha_{1}\alpha_{\mu}\rangle\right),$ (83)
@@ III.3 Gapless excitations of the parent Hamiltonian
Having established that the parent Hamiltonian is equivalent to the SU($N$) ferromagnetic Heisenberg model, we now turn to its low-energy excitations. It is well known that the ferromagnetic Heisenberg model supports gapless magnon excitations [58], which naturally provide candidates for the AQMBS states in our framework. In the following, we first impose periodic boundary conditions and construct one-magnon excitations with well-defined momentum. Subsequently, we analyze the case of open boundary conditions, where the corresponding low-energy excitations are obtained by constructing standing-wave-like states. Similar calculations have been performed in Appendix C of Ref. [33].
@@
with $p=0,1,\ldots,L-1$ and $\mathcal{N}_{p}$ being a normalization constant. Using these eigenmodes, we define a standing-wave creation operator
@@
so that $\hat{F}^{\mu 0}_{\mathrm{OBC}}(p)|\bm{0}\rangle$ is a normalized one-magnon state. Direct calculations show that
@@
Using the commutation relation $\hat{H}_{\rm p},\hat{F}_{\rm tot}^{\mu 0}]=0$ and the fact that $|\bm{0}\rangle$ is a ground state of the SU($N$) ferromagnetic Heisenberg model, we can show that the following states are also ground states of the parent Hamiltonian: